# Gamma positivity, PL homeomorphism types, and orthogonal polynomials

### 23 Mar 2026 v2

@@ 2.2.2 Behavior in large dimensions
Since $\Delta$ is a simplicial sphere in our setting, the $M$-vector condition on $g(\Delta)$ (Part 2c) of Theorem 1.6) and Corollary 2.8 imply that $g_{k+2}(\Delta)\leq g_{k+1}(\Delta)^{\langle k\rangle}\leq g_{k+1}(\Delta)^{\frac{k+2}{k+1}}$. This means that
@@
Let $\Delta$ be a (d1)(d-1)-dimensional simplicial sphere satisfying the link condition such that $\gamma(\Delta)\geq 0$. Suppose that $d=o(g_{1}(\Delta))$ or $d^{2}=o(g_{2}(\Delta))$ with respect to $d$. Note that the latter implies the former. Fix $1\leq k\leq\frac{d}{2}-2$.
@@
    This is implied by our discussion above since the remaining terms of $\frac{\mathcal{C}_{k}}{g_{k+1}(\Delta)}$ are either quadratic polynomials in $d$ or bounded above and below by constant multiples of them. Note that 2k(dk+1)2k(d-k+1) with $0\leq k\leq\frac{d}{2}$ is also bounded above by a constant multiple of a quadratic polynomial in $d$ when $k$ is a sublinear function in $d$ (where we can also phrase a more precise version).
@@
    $g_{k+2}(\Delta)\leq g_{k+1}(\Delta)^{\langle k+1\rangle}\leq g_{k+1}(\Delta)^{\frac{k+1}{k}}\Longrightarrow\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)^{\frac{k+1}{k}}}\leq 1\leq f_{1}(\Delta)^{\frac{1}{k}}.$
@@
    with respect to $d$, then the denominator may possibly have a limit in $(0,1)$ as $d\to+\infty$ while the numerator of the fraction in the upper bound is 1. However, decreasing the denominator while increasing the numerator would give rise to an even larger upper bound which is “more trivial” than the one above. Thus, it suffices to only assume that the asymptotic upper bound for $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$.
@@
Unwinding the definitions of $\mathcal{C}_{k}$ and $\mathcal{C}_{k-1}$, the inequality above can be interpreted as a lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ in terms of $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$, $g_{k}(\Delta)$, and $f_{1}(\Delta)$. Note that $f_{1}(\Delta)^{\frac{1}{k}}\geq 1$ and $g_{k}(\Delta)^{\frac{1}{k}}\geq 1$ for each $k$ since $f_{1}(\Delta)>\binom{d+1}{2}>1$ and $g_{k}(\Delta)\geq\binom{d}{k}-\binom{d}{k-1}\geq 1$ for $0\leq k\leq\frac{d}{2}$ if $\gamma(\Delta)\geq 0$. We also have $\lim_{k\to+\infty}f_{1}(\Delta)^{\frac{1}{k}}=1$. Note that $f_{1}(\Delta)$ does not depend on $k$. We can describe how the induced lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ increases/decreases with respect to these parameters.
  • Keeping $g_{k}(\Delta)$ and $f_{1}(\Delta)$ fixed, we see that increasing $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$ decreases the numerator and increases the denominator of the upper bound of $1-\frac{1}{f_{1}(\Delta)}\frac{\mathcal{C}_{k}}{g_{k+1}(\Delta)}$. Since this gives decreases the upper bound of $1-\frac{1}{f_{1}(\Delta)}\frac{\mathcal{C}_{k}}{g_{k+1}(\Delta)}$, we have a larger lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$.
@@
$\lim_{d\to+\infty}f_{1}(\Delta)^{\frac{1}{k}}=+\infty,$
it may still be possible that
@@
Let $\Delta$ be a (d1)(d-1)-dimensional simplicial sphere satisfying the link condition. Consider the inequality
@@
    Suppose that $1\leq k\leq\frac{d}{2}-1$ is a positive constant or $\lim_{d\to+\infty}k=+\infty$. If $\lim_{d\to+\infty}f_{1}(\Delta)^{\frac{1}{k}}=+\infty$ (e.g. $k$ a constant or the small functions in Proposition 2.44), the limit of the upper bound of the inequality as $d\to+\infty$ is equal to 0 only if
@@
    Since $\lim_{d\to+\infty}k=+\infty$ and
@@
      If $k$ is a constant, this is not possible.
@@
      If $k$ is a constant, this is not possible.
@@
      Since $d=o(g_{1}(\Delta))$, the terms $dg_{1}(\Delta)$ is asymptotically strictly larger than $\binom{d+1}{2}$, which is a quadratic polynomial in $d$, which is a quadratic polynomial in $d$. The term $g_{2}(\Delta)$ is the only other one that could have this property. This implies that
@@
      If $g_{1}(\Delta)$ is linear in $d$, then $dg_{1}(\Delta)$ and $g_{2}(\Delta)$ are both quadratic in $d$ (Proposition 2.11). Suppose that $g_{1}(\Delta)$ is linear in $d$ with $\lim_{d\to+\infty}\frac{k}{d}=\alpha$ for some $0<\alpha\leq\frac{1}{2}$. Then, dividing the numerator and denominator of the limit by $d^{2}$ and taking limits gives the claimed limit.
@@
      The case where $k$ is a constant follows from dividing the numerator and denominator by $d^{2}$ since each term added to form $\frac{\mathcal{C}_{k-1}}{g_{k}(\Delta)}$ is at most linear in $d$.
@@ 2.2.3 Summary and asymptotic behavior of individual inequalities
Expanding the averaged terms $\mathcal{C}_{k}$ and $\mathcal{C}_{k-1}$ in terms of global parts, this gives rise to a lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ in terms of $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$, $f_{1}(\Delta)$, and $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)^{\frac{k+1}{k}}}$. The simplicial sphere $\Delta$ having a smaller number of edges or the $M$-vector inequality on $g(\Delta)$ (Part 2c) of Theorem 1.6) would yield a tighter bound (i.e. larger lower bound). When $d\gg 0$, the lower bound increases as $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$ increases.
@@
In general, we take a closer look at the specific types of bounds obtained when $\dim\Delta$ is large. This involves behavior of “limiting inequalities” as $d\to+\infty$. We have considered index parameters $1\leq k\leq\frac{d}{2}$ that are be linear or sublinear (e.g. constant) in $d$ (e.g. see Proposition 2.40 and Corollary 2.42). In this section, we focus on the case where $k$ is constant. It describes limiting behavior of individual link condition inequalities, while the other types of functions describe those of collections of link inequalities. This is where we will mainly start using the assumption $\gamma(\Delta)\geq 0$. The induced limit inequality (i.e. for $d\gg 0$) holds trivially if there is a constant $C>0$ such that
@@
for (finite) constant $\beta>0$.
@@
    The upper bound $\sigma$ in the induced limiting link condition inequality $\sigma\geq 1-\beta$ as $d\to+\infty$ can be interpreted asymptotically as the quotient of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ by $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$ (Part 1 and Part 3 of Corollary 2.48). In the context of Remark 2.43, this would be a constraint on the rate of decrease of the terms $\frac{g_{m+1}(\Delta)}{g_{m}(\Delta)}$.
@@
    While the lower bound for the quotient of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ by $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$ from Corollary 2.33 can be translated to one for $\gamma_{k+2}(\Delta)$, the gamma vector components $\gamma_{i}(\Delta)$ affected are mainly those that are asymptotically dominant in $g_{i}(\Delta)$. For example, having $\gamma_{\ell+1}(\Delta)\geq(d+1)g_{\ell}(\Delta)$ (e.g. when $g_{\ell+1}(\Delta)\geq 2(d+1)g_{\ell}(\Delta)$) would imply that $\gamma_{\ell+1}(\Delta)$ must be the asymptotically dominant term in $g_{\ell+1}(\Delta)$ (Corollary 2.4). If this holds for all $\ell\in\{k,k+1,k+2\}$, we can take the inequality to be an asymptotic lower bound on the quotient of $\frac{\gamma_{k+2}(\Delta)}{\gamma_{k+1}(\Delta)}$ by $\frac{\gamma_{k+1}}{\gamma_{k}}$.

### 26 Mar 2026 v3

@@ 2.2.2 Behavior in large dimensions
Since $\Delta$ is a simplicial sphere in our setting, the $M$-vector condition on $g(\Delta)$ (Part 2c) of Theorem 1.7) and Corollary 2.8 imply that $g_{k+2}(\Delta)\leq g_{k+1}(\Delta)^{\langle k+1\rangle}\leq g_{k+1}(\Delta)^{\frac{k+2}{k+1}}$. This means that
@@
Let $\Delta$ be a (d1)(d-1)-dimensional simplicial sphere satisfying the link condition such that $\gamma(\Delta)\geq 0$. Suppose that $d=o(g_{1}(\Delta))$ with respect to $d$ (e.g. when $d^{2}=o(g_{2}(\Delta))$). Fix $1\leq k\leq\frac{d}{2}-2$.
@@
    This is implied by our discussion above since the remaining terms of $\frac{\mathcal{C}_{k}}{g_{k+1}(\Delta)}$ are either quadratic polynomials in $d$ or bounded above and below by constant multiples of them. Note that 2k(dk+1)2k(d-k+1) with $0\leq k\leq\frac{d}{2}$ is also bounded above by a constant multiple of a quadratic polynomial in $d$ when $k$ is a sublinear function in $d$ (where we can also phrase a more precise version). The comparison between the examples follows from $g_{2}(\Delta)\leq g_{1}(\Delta)^{\langle 1\rangle}=\binom{g_{1}(\Delta)+1}{2}$.
@@
    $g_{k+1}(\Delta)\leq g_{k}(\Delta)^{\langle k\rangle}\leq g_{k}(\Delta)^{\frac{k+1}{k}}\Longrightarrow\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)^{\frac{k+1}{k}}}\leq 1\leq f_{1}(\Delta)^{\frac{1}{k}}.$
@@
    with respect to $d$, then the denominator may have a value in $(0,1)$ as $d\to+\infty$ while the numerator of the fraction in the upper bound is 1. However, decreasing the denominator while increasing the numerator would give rise to an even larger upper bound which is “more trivial” than the one above. Thus, it suffices to only assume that the asymptotic upper bound for $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$.
@@
Unwinding the definitions of $\mathcal{C}_{k}$ and $\mathcal{C}_{k-1}$, the inequality above can be interpreted as a lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ in terms of $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$, $g_{k}(\Delta)$, and $f_{1}(\Delta)$. Note that $f_{1}(\Delta)^{\frac{1}{k}}\geq 1$ and $g_{k}(\Delta)^{\frac{1}{k}}\geq 1$ for each $k$ since $f_{1}(\Delta)>\binom{d+1}{2}>1$ and $g_{k}(\Delta)\geq\binom{d}{k}-\binom{d}{k-1}\geq 1$ for $0\leq k\leq\frac{d}{2}$ if $\gamma(\Delta)\geq 0$. Given a fixed $d$, we have $\lim_{k\to+\infty}f_{1}(\Delta)^{\frac{1}{k}}=1$. Note that $f_{1}(\Delta)$ does not depend on $k$. We can describe how the induced lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ increases/decreases with respect to these parameters.
  • Keeping $g_{k}(\Delta)$ and $f_{1}(\Delta)$ fixed, we see that increasing $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$ decreases the numerator and increases the denominator of the upper bound of $1-\frac{1}{f_{1}(\Delta)}\frac{\mathcal{C}_{k}}{g_{k+1}(\Delta)}$. Since this decreases the upper bound of $1-\frac{1}{f_{1}(\Delta)}\frac{\mathcal{C}_{k}}{g_{k+1}(\Delta)}$, we have a larger lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$.
@@
$\lim_{d\to+\infty}f_{1}(\Delta)^{\frac{1}{k}}=+\infty$
for each fixed $k$, it may still be possible that
@@
Let $\Delta$ be a (d1)(d-1)-dimensional simplicial sphere satisfying the link condition such that $\gamma(\Delta)\geq 0$. Consider the inequality
@@
    Suppose that $1\leq k\leq\frac{d}{2}-1$ is a positive constant or $\lim_{d\to+\infty}k=+\infty$. If $\lim_{d\to+\infty}f_{1}(\Delta)^{\frac{1}{k}}=+\infty$ (e.g. $k$ a constant or the small functions in $d$ from Proposition 2.44), the limit of the upper bound of the inequality as $d\to+\infty$ is equal to 0 only if
@@
    Since $\lim_{d\to+\infty}k=+\infty$, $\lim_{d\to+\infty}f_{1}(\Delta)^{\frac{1}{k}}=+\infty$, and
@@
      If $k$ is a constant, this is impossible.
@@
      If $k$ is a constant, this is impossible.
@@
      Since $d=o(g_{1}(\Delta))$, the terms $dg_{1}(\Delta)$ is asymptotically strictly larger than $\binom{d+1}{2}$, which is a quadratic polynomial in $d$. The term $g_{2}(\Delta)$ is the only other one that could have this property. This implies that
@@
      $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}=o\left(\frac{g_{2}(\Delta)}{d^{2}}\right).$
@@
      If $g_{1}(\Delta)$ is linear in $d$, then $dg_{1}(\Delta)$ and $g_{2}(\Delta)$ are both quadratic in $d$ (Proposition 2.11). Suppose that $k$ is linear in $d$ with $\lim_{d\to+\infty}\frac{k}{d}=\alpha$ for some $0<\alpha\leq\frac{1}{2}$. Then, dividing the numerator and denominator of the limit by $d^{2}$ and taking limits gives the claimed limit.
@@
      The case where $k$ is a constant follows from dividing the numerator and denominator by $d^{2}$ since each term added to form $\frac{\mathcal{C}_{k-1}}{g_{k}(\Delta)}$ is at most linear in $d$ by Proposition 2.11 (also see Section 2.2.3).
@@ 2.2.3 Summary and asymptotic behavior of individual inequalities
$1-\frac{1}{f_{1}(\Delta)}\frac{\mathcal{C}_{k}}{g_{k+1}(\Delta)}\leq f_{1}(\Delta)^{\frac{1}{k}}\frac{\left(1-\frac{1}{f_{1}(\Delta)}\frac{\mathcal{C}_{k-1}}{g_{k}(\Delta)}\right)^{\frac{k+1}{k}}}{\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)^{\frac{k+1}{k}}}}=f_{1}(\Delta)^{\frac{1}{k}}\frac{\left(1-\frac{1}{f_{1}(\Delta)}\frac{\mathcal{C}_{k-1}}{g_{k}(\Delta)}\right)^{\frac{k+1}{k}}}{\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}}\cdot g_{k}(\Delta)^{\frac{1}{k}}$
@@
Expanding the averaged terms $\mathcal{C}_{k}$ and $\mathcal{C}_{k-1}$ in terms of global parts, this gives rise to a lower bound of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ in terms of $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$, $f_{1}(\Delta)$, and $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)^{\frac{k+1}{k}}}$. The simplicial sphere $\Delta$ having a smaller number of edges or the $M$-vector inequality on $g(\Delta)$ (Part 2c) of Theorem 1.7) would yield a tighter bound (i.e. larger lower bound). When $d\gg 0$, the lower bound increases as $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$ decreases.
@@
In general, we take a closer look at the specific types of bounds obtained when $\dim\Delta$ is large. This involves behavior of “limiting inequalities” as $d\to+\infty$. We have considered index parameters $1\leq k\leq\frac{d}{2}$ that are be linear or sublinear (e.g. constant) in $d$ (e.g. see Proposition 2.40 and Corollary 2.42). In this section, we focus on the case where $k$ is constant. It describes limiting behavior of individual link condition inequalities, while the other types of functions describe those of collections of link condition inequalities. This is where we will mainly start using the assumption $\gamma(\Delta)\geq 0$. The induced limit inequality (i.e. for $d\gg 0$) holds trivially if there is a constant $C>0$ such that
@@
for (finite) constant $\beta\geq 0$.
@@
    The upper bound $\sigma$ in the induced limiting link condition inequality $\sigma\geq 1-\beta$ as $d\to+\infty$ can be interpreted asymptotically (with respect to $d$) as the quotient of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ by $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$ since we are dividing by the constant $k+2$ (Part 1 and Part 3 of Corollary 2.48). In the context of Remark 2.43, this would be a constraint on the rate of decrease of the terms $\frac{g_{m+1}(\Delta)}{g_{m}(\Delta)}$.
@@
    While the lower bound for the quotient of $\frac{g_{k+2}(\Delta)}{g_{k+1}(\Delta)}$ by $\frac{g_{k+1}(\Delta)}{g_{k}(\Delta)}$ from Corollary 2.33 can be translated to one for $\gamma_{k+2}(\Delta)$, the gamma vector components $\gamma_{i}(\Delta)$ affected are mainly those that are asymptotically dominant in $g_{i}(\Delta)$. For example, having $\gamma_{\ell+1}(\Delta)\geq(d+1)g_{\ell}(\Delta)$ (e.g. when $g_{\ell+1}(\Delta)\geq 2(d+1)g_{\ell}(\Delta)$) would imply that $\gamma_{\ell+1}(\Delta)$ must be the asymptotically dominant term in $g_{\ell+1}(\Delta)$ (Corollary 2.4). If this holds for all $\ell\in\{k,k+1,k+2\}$, we can take the inequality to be an asymptotic lower bound on the quotient of $\frac{\gamma_{k+2}(\Delta)}{\gamma_{k+1}(\Delta)}$ by $\frac{\gamma_{k+1}(\Delta)}{\gamma_{k}(\Delta)}$.